OpenStudy (chillhill):

How do you make x the subject: ax-3y=b(x+2y)

OpenStudy (imtiaz7):

you mean that to separate the x?

OpenStudy (chillhill):

yeah

OpenStudy (imtiaz7):

so ax-3y=b(x+2y) ax-3y=bx+2by ax-bx=2by+3y x(a-b)=y(2b+3) x=y(2b+3)/(a-b)

OpenStudy (chillhill):

thank you

OpenStudy (imtiaz7):

welcome :) @chillhill